What is the midpoint between (-8, 5) and (2, -2)?


(-5, 1.5)
(-5, 3.5)
(-6, 3)
(-3, 1.5)

Answer by jim_thompson5910(35256)

Average the x coordinates: (-8+2)/2 = -6/2 = -3


Average the y coordinates: (5+(-2))/2 = 3/2 = 1.5


So the midpoint is (-3, 1.5)